# Stereographic Projections for Designs on the Sphere

### v1

@@ Abstract
This paper is concerned with the use of the stereographic projection in order to map the sphere points of designs for regression models in three variables and design space the unit ball onto the two-dimensional stereogram. Details of the projection and its attendant stereogram are given. Stereograms which represent the spherical component of designs with points on the sphere and the centre of the unit ball, including the central composite and the Box-Behnken designs, are introduced. In addition, an example in which stereograms are used to elucidate the geometric isomorphism of a suite of designs generated by an exchange algorithm is presented.
keywords:
Stereographic projection; stereogram; quadratic polynomial regression model; unit ball; geometric isomorphism1 Introduction
Designs over the unit ball which comprise a finite set of support points have been reported sporadically in the literature as, for example, in the papers of Bose and Draper (1959), Hardin and Sloane (1991), Gilmour (2006), Dette and Grigoriev (2014), Hirao et al. (2015) and Radloff and Schwabe (2023). The designs are not straightforward to construct and, to compound matters, the three-dimensional prototypes, which help to convey the innate structure of the designs, are not easy to represent on paper. The aim of the present study therefore is to introduce the stereographic projection as a valuable tool, in research and pedagogically, for representing the spherical components of designs in three independent variables over the unit ball in two dimensions. The paper is organized as follows. The notion of the stereographic projection and the construction of the attendant stereogram are presented in Section 2. Illustrative examples relating to the quadratic polynomial model in three variables are then provided in Section 3. Some general comments on the relevance of the use of the stereographic projection in experimental design are presented in Section 4.
2 The Stereographic Projection
@@
Consider first a perfect terrestrial globe which is centred at the origin with the North-South axis oriented vertically and the equatorial plane perpendicular to that axis. The stereographic projection maps points and circles on the three-dimensional globe onto the two-dimensional equatorial plane, with the resulting construct termed a stereogram. In particular, consider a point lying on the surface of the globe in the northern hemisphere. Then a line can be drawn from that point to the South Pole and the point of intersection of the line with the equatorial plane indicated on the stereogram with a solid circle, as shown in Figure 1(a). Similarly, a line can be drawn from a point on the surface of the globe in the southern hemisphere to the North Pole and its intersection with the equatorial plane denoted for clarity by an open circle, as in Figure 1(b). In addition, parallels, that is lines of latitude, are projected onto the equatorial plane as circles centred at the origin and meridians, that is lines of longitude, as straight lines through the origin. These constructs can be represented by a polar net, as shown in Figure 1(c). Great circles through two points on the surface of the sphere can also be projected onto the equatorial plane, specifically as arcs with diagonally opposite end points touching the equator. Selected meridians and great circles are shown on the stereogram in Figure 1(d).
Three-dimensional directions are often indicated on the stereogram and, for example, the vectors and are denoted compactly by and . To give a sense of the orientation of the equatorial plane, selected directions are included in Figure 1(d). Finally, note that a small circle drawn on the surface of the globe can be projected as a circle onto the stereogram but the centres of the two circles do not necessarily correspond. Further details on the stereographic projection are given, for example, in the books on crystallography by Borchardt-Ott (2011), De Graef and McHenry (2012) and Hammond (2015).
@@
The stereographic projection can now be adapted to accommodate designs in three dimensions on design space the unit ball with centre at the origin and support points on the boundary and at the centre of the ball. Specifically, consider the perfect terrestrial globe described above as a unit sphere with a right-handed coordinate system specified by the coordinates and . Then the point can be taken taken to coincide with the North Pole, the point with the intersection of a line in the direction with the equator and the -plane with the equatorial plane. The support points of a design on the surface of the sphere can be projected onto the equatorial plane, that is the stereogram, by invoking the mapping . Projections onto the stereogram of the appropriate meridians and great circles which connect the design points then follow.
3 Examples
3.1 Subset Designs
@@
To fix ideas, consider the subset designs for the quadratic polynomial model with three independent variables over the unit ball ntroduced into the literature in the paper by Gilmour (2006). The support points of the designs lie on the surface and at the centre of the unit ball and are taken from the sets specified by
@@
which comprise and points respectively. The subscript ‘cycle’ indicates that the three coordinates of the points can be moved cyclically. More specifically, subset designs can be expressed as a linear combination of these sets of points, that is as , where and the remaining coefficients, and , are greater than or equal to zero and are chosen in such a way that the moment matrix is nonsingular. From a symmetry perspective, the sets of points and correspond to the centre point and, for a cube inscribed in the unit sphere, to the intersection of the six normals to the faces of the cube with the sphere, to the intersection of the 12 lines from the origin through the mid-points of the edges of the cube with the sphere, and to the eight vertices of the cube itself, respectively. It thus follows that the points on the unit sphere of a subset design with equal weights assigned to the subsets and belong to the point group .
@@
The central composite design over design space the unit ball is the subset design and the stereographic projection of the 14 design points on the sphere yields the stereogram shown in Figure 2(a). Similarly, the Box-Behnken design is the subset design and the stereogram of the 12 points on the sphere is shown in Figure 2(b). Note that the points in and are coloured in the stereograms in black, blue and red, respectively, in Figure 2 and that the complementary relationship between the two designs is clearly illustrated by the stereograms. Subset designs with different coefficients allocated to the sets of points and can also be represented compactly on a stereogram by using colour-coding and points with areas proportional to their relative weightings. For example, the subset design is rotatable and the different allocations of points to the two sets and is highlighted in the stereogram in Figure 2(c). In addition, the approximate -optimal design for the quadratic polynomial model in three variables on the unit ball derived by Dette and Grigoriev (2014) is the subset design and thus a weighted form of the central composite design. The stereogram of the sphere points of the design and their relative weighting is shown in Figure 2(d). Note that the design itself is rather curious in that more than half of the weight is placed on the centre point.
3.2 An Application
@@
Gilmour and Trinca (2012), in Example 2 of their paper, considered optimal designs which accommodate inference for a full quadratic polynomial model in three independent variables with support points on the surface and at the centre of a ball with radius 3. For clarity, their results are reported here for the unit ball. The authors focussed on nonsingular designs with 18 support points taken, with replacement, from the individual points belonging the subsets and of the previous example. More specifically, they considered, inter alia, D-optimal designs which minimise the volume of the confidence region for the parameters of the model with the intercept treated as a nuisance parameter and (DP)-optimal designs which minimise the D-criterion multiplied by an appropriate quantile of the -distribution raised to the power equal to the number of parameters of interest, in this case nine. It can in fact be easily shown that over possible designs must be enumerated in order to identify those designs which are strictly D- and (DP)-optimal. As a consequence, Gilmour and Trinca (2012) used a heuristic, the exchange algorithm documented in, for example, Atkinson et al. (2007), to search for designs which are optimal or near-optimal.
@@
The question now arises as to how many strictly D- and (DP)-optimal designs are generated by the exchange algorithm and, if more than one, how the resultant designs are related. This question in turn impacts on the performance of the algorithm. To this end, the exchange procedure of Gilmour and Trinca (2012) was run five million times in the programming language gauss (2022) and the requisite optimal designs which were distinct, that is non-coincident, identified. Eight distinct D-optimal designs with two centre points and distinct (DP)-optimal designs, with including a single centre point and two centre points, were so generated. The designs were then examined in the context of isomorphism and, since the factors and are quantitative, in that of geometric isomorphism (Cheng and Ye 2004). Specifically, two designs are said to be geometrically isomorphic if the designs can be obtained, one from the other, by rotations and reflections within the design space. In the present example, the isomorphic properties of the designs are determined solely by the spherical points and the actions of rotation and reflection can therefore be mapped, albeit indirectly, onto the stereograms. It is crucial to observe here that a design which is geometrically isomorphic to a D-optimal or a (DP)-optimal design is itself necessarily D-optimal or (DP)-optimal.
@@
The D-optimal design given in Table 2 of Gilmour and Trinca (2012) comprises 16 points on the sphere, none of which are replicated, and two centre points, and coincides with one of the eight distinct D-optimal designs generated by the exchange algorithm. The stereogram of the spherical points of the design is presented in Figure 3(a), with the great circles introduced as guidelines for the placement of the points. It is immediately clear from the figure that a further three distinct designs can be generated from the design by repeated rotations through about the axis and the associated stereograms are shown in Figures 3(b), 3(c) and 3(d). A further four designs can then be obtained by reflecting the designs shown in Figure 3 through the plane perpendicular to the axis, that is the -plane. The resultant eight designs so obtained are geometrically isomorphic and therefore D-optimal. These isomorphic designs therefore account for the eight distinct D-optimal designs generated by the exchange algorithm.
@@
The problem of identifying (DP)-optimal designs which are geometrically isomorphic from the 1,080 optimal designs generated by the exchange algorithm is a little more nuanced than that relating to the D-optimal designs. Consider first the design reported in Table 2 of Gilmour and Trinca (2012) which is easily identified as one of the 216 distinct (DP)-optimal designs with a single centre point generated by the exchange algorithm.
This -point design is based on one centre point and on a set of nine distinct points on the sphere, specifically the set
termed here the basis set, with each point replicated twice except for the point which occurs once. A stereogram of the 17 points on the sphere of the design is presented in Figure 4(a), with the point occurring once indicated in red.
@@
Two distinct designs can be obtained by rotation twice through about the axis defined by the vector and are geometrically isomorphic to the original design of Gilmour and Trinca (2012) and therefore (DP)-optimal. To illustrate, the stereograms of the two resultant designs are shown in Figures 4(b) and 4(c). As an aside, the designs can also be constructed by cycling the original axes of Gilmour and Trinca (2012) twice as and . Eight geometrically isomorphic designs can now be generated from each of the three designs displayed in Figures 4(a), (b) and (c) by rotation through about the -axis and reflection about an appropriate plane to yield a total of designs. The design of Gilmour and Trinca (2012) is therefore representative of a class of 24 geometrically isomorphic designs.
@@
A cursory examination of the results from the exchange algorithm indicated that designs with one centre point and any one of the nine points in the basis set occurring once and not twice are (DP)-optimal. In terms of geometric isomorphism, consider now a design constructed directly from that of Gilmour and Trinca (2012) but with the the point now replicated twice and the point occurring once. For clarity, a stereogram of the design is shown in Figure 4(d), with the point occurring once indicated in red. It is clear that the design so constructed is not geometrically isomorphic to the design of Gilmour and Trinca (2012). However, by using arguments based on rotations and reflections, it is straightforward to show that the design represents a class of 24 isomorphic (DP)-optimal designs. Furthermore, the same result holds for all designs constructed with one of the points in the set occurring once. There are therefore nine classes of designs, each comprising 24 geometrically isomorphic designs, and thus a total of designs. This result is in accord with the fact that (DP)-optimal designs with one centre point were identified from multiple runs of the exchange algorithm.
@@
Consider now the (DP)-optimal designs with two centre points generated from multiple runs of the exchange algorithm. A cursory examination of the results indicated that the points on the sphere of the design comprise two distinct points from the basis set chosen to occur once and the seven remaining points in replicated twice. The designs so defined were then constructed independently and shown to be (DP)-optimal. It now follows from arguments similar to those presented for the (DP)-optimal designs with one centre point that each of the designs so constructed represents a class of geometrically isomorphic designs, thereby yielding designs which are (DP)-optimal and accounting for the designs generated by the exchange algorithm.
4 Discussion
The focus of the present study is on the use of the stereographic projection within the context of experimental design in order to map points on the three-dimensional unit sphere onto the stereogram in two dimensions. The stereogram gives a clear visualisation of the spatial arrangement of the points and can be manipulated in order to emulate rotations of, and reflections about, the main axes. Stereograms for subset designs for the quadratic polynomial model in three independent variables over the unit ball, including the central composite and Box-Behnken designs, are presented in order to fix ideas. An example of the effective use of the stereogram is then introduced and is based on a suite of designs generated numerically by an exchange algorithm. Specifically, it is shown that classes of designs which are geometrically isomorphic can be identified by a judicious choice of operations on the stereogram. More broadly, generating optimal designs by invoking a heuristic invariably yields large numbers of similar designs and the present example reveals how such designs can be inter-related and, thereby, classified.

### v2

@@ Abstract
This paper is concerned with the use of the stereographic projection to map the points of a design on the sphere in three dimensions onto a two-dimensional stereogram. Details of the projection and its attendant stereogram are given and the application of the constructs to designs on the sphere is highlighted. Spherical -designs are then introduced and stereograms of selected examples are presented. In addition, examples of regression models with design space the unit ball are considered, stereograms which represent the spherical component of such designs are given and an example in which stereograms are used to elucidate the geometric isomorphism of a suite of designs generated by an exchange algorithm is presented. Emphasis throughout the text is placed on the use of the stereographic projection in teaching and research within the framework of statistics.
Keywords: Stereographic projection; stereogram; spherical -designs; regression models; geometric isomorphism
1. Introduction
The aim of the present study is to introduce the stereographic projection as a valuable tool in teaching and research for representing the points of a design over the three-dimensional sphere as a stereogram in two dimensions. The projection is used extensively in, for example, crystallography, complex analysis and geology but has received scant attention in the statistics and related literature. The designs of interest here arise in areas as diverse as the distribution of charged particles on the sphere, the nature and application of spherical -designs and regression models with design space the unit sphere.
The paper is organized as follows. The notion of the stereographic projection and the construction of the attendant stereogram are presented in Section 2. Spherical -designs are introduced in Section 3 and stereograms of selected examples are presented there. In Section 4 an example of a quadratic regression model in three variables, taken from the statistical literature, is discussed and the use of stereograms to elucidate the geometric isomorphism of a suite of designs generated from the model by an exchange algorithm is presented. Some general comments are included in Section 5. The Schönflies notation for points groups is used to describe the structures of the designs of interest formally. An R package for constructing the stereograms of the designs is also available from the author on request.
2. The Stereographic Projection
@@
Consider first a perfect terrestrial globe which is centred at the origin with the North-South axis oriented vertically and the equatorial plane perpendicular to that axis. The stereographic projection maps points and circles on the three-dimensional globe onto the two-dimensional equatorial plane, with the resulting construct termed a stereogram. In particular, consider a point lying on the surface of the globe in the northern hemisphere. Then a line can be drawn from that point to the South Pole and the point of intersection of the line with the equatorial plane indicated on the stereogram with a solid circle, as shown in Figure 1(a). Similarly, a line can be drawn from a point on the surface of the globe in the southern hemisphere to the North Pole and its intersection with the equatorial plane denoted, for clarity, by an open circle, as in Figure 1(b). In addition, parallels, that is lines of latitude, are projected onto the equatorial plane as circles centred at the origin and meridians, that is lines of longitude, as straight lines through the origin. These constructs can be represented by a polar net, as shown in Figure 1(c). Great circles through two points on the surface of the sphere can also be projected onto the equatorial plane as arcs with diagonally opposite end points touching the equator. Selected meridians and great circles are shown on the stereogram in Figure 1(d). Three-dimensional directions are often indicated on the stereogram and, for example, the vectors and are denoted compactly by and . To give a sense of the orientation of the equatorial plane, selected directions are included in Figure 1(d). Finally, note that a small circle drawn on the surface of the globe or sphere can be projected onto the stereogram as a circle but the centres of the two circles do not necessarily correspond. Note also that the stereographic projection preserves angles but not distances on the sphere. A point group in three dimensions, that is a set of symmetry operations which leave at least one point unchanged, forms an integral part of the stereographic projection, with the unchanged point taken to lie at the centre of the sphere. In addition, orbits are very often used to highlight the symmetries which define a point group. Specifically, an orbit is a set of points which can be obtained by a group action of the point group on a given point in that set. Point groups are presented in the Schönflies notation throughout the text. Further details on the stereographic projection are given, for example, in the books on crystallography by Borchardt-Ott (2011), De Graef and McHenry (2012) and Hammond (2015). Finally, it is worth noting that the exponential map, which can also be invoked to represent points on a sphere in two dimensions, preserves distances but not angles and is not relevant in the present context.
@@
The stereographic projection can immediately be adapted to accommodate designs in three dimensions over the unit sphere. Specifically, consider the perfect terrestrial globe described above as a unit sphere with a right-handed coordinate system specified by the coordinates and . Then the point can be taken taken to coincide with the North Pole, the point with the intersection of a line in the direction with the equator and the -plane with the equatorial plane. The support points of a design on the surface of the sphere can be projected onto the equatorial plane, that is the stereogram, by invoking the mapping . Projections onto the stereogram of the appropriate meridians and great circles which connect the design points then follow.
3. Spherical -Designs
Spherical -designs were introduced into the literature on algebraic combinatorics by Delsarte et al. (1977) and are of considerable current interest both theoretically and practically. In the present context of three dimensions, the designs comprise points on the sphere for which the average evaluated at a homogeneous spherical polynomial of degree is exactly equal to the integral of the polynomial with respect to the uniform distribution on the sphere. More formally, a spherical -design is a finite subset on the unit sphere which satisfies the cubature relationship
where is a uniform measure on the sphere and denotes the space of homogeneous spherical polynomials of degree less than or equal to . This definition can be modified, following Gräf and Potts (2011) and Sawa et al. (2019, p. 14), by replacing the sum on the right-hand of Equation (1) with a weighted sum , where represents a weight associated with such that . The resultant designs are referred to in the present context as weighted spherical -designs.
A number of examples of spherical -designs and their weighted counterparts are now introduced in order to highlight the use of the stereogram in identifying the nature and structure of the designs. Spherical -designs in three dimensions are, in general, not straightforward to construct and the computational problems are highlighted in the papers by Hardin and Sloane (1996) and Gräf and Potts (2011). The examples presented here are drawn in digital form from Neil Sloane’s website Sloane (1996) for the spherical -designs and from the website of Gräf (2011) for the weighted spherical -designs.
Example 3.1: Spherical -designs: Stereograms of - and -point spherical 5-designs and 60-point spherical -designs with and are given in Figures 2(a), (b), (c) and (d), respectively, with the appropriate point groups represented in Schönflies notation in the captions. The -point and -point designs have the same values but very different structures. Specifically, it is clear from the stereograms that the -point designs has tetrahedral symmetry, T, while the -point design has a five-fold principal axis and symmetry D5d. The two -point designs have the same number of points but values of and and are structurally very different. Thus, as shown in stereogram 2(c), the -point design with has icosahedral symmetry with point group Ih and represents a football with hexagonal and pentagonal faces connected by edges. As an aside, it also represents the carbon allotrope, the Buckminsterfullerene, and its semi-spheres are geodesic domes. In contrast, the -point spherical -design shown in stereogram 2(d) comprises five snub tetrahedra and exhibits tetrahedral symmetry with point group T.
Example 3.2: Weighted spherical -designs
Stereograms of the 22-point spherical 5-design and the weighted 22-point spherical -design are shown in Figures 3(a) and 3(b). The two designs clearly belong to the same point group, that is D5d, and comprise orbits defined by the North and South Poles and by points located on two sets of north and south latitudes which are at the same polar angle from the equator. The designs do however exhibit subtle differences. First, it is easily shown that the orbits comprising points located on the latitudes of the two designs are defined by different polar angles and that the latitudes associated with the weighted spherical -design are closer to the equator. Second, and more importantly, the spherical 5-design comprises equally weighted points while, by definition, the weighted spherical -design does not. Specifically, the orbits of the latter design are indicated in black, red and blue in Figure 3(b) and have weights and , respectively, that is taken from the centre outwards. Stereograms of the 28-point spherical 6-design and its counterpart, the weighted 28-point spherical 8-design, are given in Figures 3(c) and 3(d). In this case, the two designs exhibit different symmetries, with the spherical -design belonging to the point group D2d and the spherical -design to the point group T. The orbits of the latter design are defined by the four points corresponding to the vertices of the tetrahedron and two sets of points located around the three-fold axes. The orbits are highlighted on the stereogram in Figure 3(d) in black, red and blue and have weights and , that is from the vertices of the tetrahedron outwards, respectively.
4. Regression Modelling
Designs over the unit ball which comprise a finite set of support points located at the centre of the ball and on the sphere have been reported sporadically in the literature as, for example, in the papers of Bose and Draper (1959), Hardin and Sloane (1991), Gilmour (2006), Dette and Grigoriev (2014), Hirao et al. (2015) and Radloff and Schwabe (2023). Three-dimensional prototypes of these designs help to convey the innate structure of the hyper-spherical designs and can be represented to some advantage by invoking the stereographic projection and attendant stereogram.
4.1 Pedagogical Examples
@@
To fix ideas, consider the subset designs for the quadratic polynomial model with three independent variables over the unit ball introduced into the literature in the paper by Gilmour (2006). The support points of the designs lie on the surface and at the centre of the unit ball and are taken from the sets specified by
@@
which comprise and points, respectively. The subscript ‘cycle’ indicates that the three coordinates of the points can be moved cyclically. More specifically, subset designs can be expressed as a linear combination of these sets of points, that is as , where and the remaining coefficients, and , are greater than or equal to zero and are chosen in such a way that the moment matrix is nonsingular. From a symmetry perspective, the sets of points and correspond to the centre point and, for a cube inscribed in the unit sphere, to the intersection of the six normals to the faces of the cube with the sphere, to the intersection of the 12 lines from the origin through the mid-points of the edges of the cube with the sphere, and to the eight vertices of the cube itself, respectively. It thus follows that the points on the unit sphere of a subset design with equal weights assigned to the subsets and belong to the point group .
@@
The central composite design over design space the unit ball is the subset design and the stereographic projection of the 14 design points on the sphere yields the stereogram shown in Figure 4(a). Similarly, the Box-Behnken design is the subset design and the stereogram of the 12 points on the sphere is shown in Figure 4(b). Note that the points in and are coloured in the stereograms in black, blue and red in Figure 4, respectively, and that the complementary relationship between the two designs is clearly illustrated by the stereograms. Subset designs with different coefficients allocated to the sets of points and can also be represented compactly on a stereogram by using colour coding and points with areas proportional to their relative weightings. For example, the subset design is rotatable and the different allocations of points to the two sets and is highlighted in the stereogram in Figure 4(c). In addition, the approximate -optimal design for the quadratic polynomial model in three variables on the unit ball derived by Dette and Grigoriev (2014) is the subset design and thus a weighted form of the central composite design. The stereogram of the sphere points of the design and their relative weighting is shown in Figure 4(d). Note that the design itself is rather curious in that more than half of the weight is placed on the centre point.
4.2 Geometric Isomorphism
@@
Gilmour and Trinca (2012), in Example 2 of their paper, considered optimal designs which accommodate inference for a full quadratic polynomial model in three independent variables with support points on the surface and at the centre of a ball. The authors focussed on nonsingular designs with 18 support points taken, with replacement, from the individual points belonging the subsets and of the previous subsection. More specifically, they considered, inter alia, DS-optimal designs which minimise the volume of the confidence region for the parameters of the model with the intercept treated as a nuisance parameter and (DP)S-optimal designs which minimise the DS-criterion multiplied by an appropriate quantile of the -distribution raised to a power equal to the number of parameters of interest, in this case nine. However, over possible designs must be enumerated in order to identify those designs which are strictly DS- and (DP)S-optimal. As a consequence, Gilmour and Trinca (2012) used a heuristic, the exchange algorithm, which is documented in, for example, Atkinson et al. (2007), to search for such designs.
@@
The question now arises as to how many strictly DS- and (DP)S-optimal designs are generated by the exchange algorithm and, if more than one, how the resultant designs are related. This question in turn impacts on the performance of the algorithm. To this end, the exchange procedure of Gilmour and Trinca (2012) was run five million times in the programming language gauss (2024) and the requisite optimal designs which were distinct identified. Eight distinct DS-optimal designs with two centre points and distinct (DP)S-optimal designs, with including a single centre point and two centre points, were so generated. The designs were then examined in the context of isomorphism and, since the factors and are quantitative, in that of geometric isomorphism (Cheng and Ye, 2004). Specifically, two designs are said to be geometrically isomorphic if the designs can be obtained, one from the other, by rotations and reflections within the design space. In the present example, the isomorphic properties of the designs are determined solely by the spherical points and the actions of rotation and reflection can therefore be mapped directly onto the appropriate stereograms.
@@
The DS-optimal design given in Table 2 of Gilmour and Trinca (2012) comprises 16 points on the sphere, none of which are replicated, and two centre points and the stereogram of the spherical points of the design is presented in Figure 5(a). It is immediately clear that a further three designs can be generated from the design by repeated anticlockwise rotations through about the axis and, to illustrate, a stereogram of one of these is shown in Figure 5(b). A further four designs can then be obtained by reflecting these designs through the plane perpendicular to the axis, that is the -plane. The resultant eight designs so obtained are geometrically isomorphic and therefore DS-optimal and account for the eight distinct DS-optimal designs generated by the exchange algorithm.
@@
The problem of identifying (DP)S-optimal designs which are geometrically isomorphic from the 1,080 optimal designs generated by the exchange algorithm is a little more nuanced than that relating to the DS-optimal designs. Consider first the design reported in Table 2 of Gilmour and Trinca (2012) which is one of the 216 (DP)S-optimal designs with a single centre point. This -point design is based on the one centre point and on the set of nine points on the sphere
@@
termed here the basis set, with each point replicated twice except for the point which occurs once. A stereogram of the 17 points on the sphere of the design is presented in Figure 6(a), with the point occurring once indicated in red. Two designs can now be obtained by rotation twice through about the axis defined by the vector and are geometrically isomorphic to the original design of Gilmour and Trinca (2012). To illustrate, the stereograms of the two resultant designs are shown in Figures 6(b) and 6(c). As an aside, the designs can also be constructed by cycling the original axes of Gilmour and Trinca (2012) twice as and . Eight geometrically isomorphic designs can now be generated from each of the three designs displayed in Figures 6(a), (b) and (c) by rotation through about the -axis and reflection about an appropriate plane to yield a total of designs. The design of Gilmour and Trinca (2012) is therefore representative of a class of 24 geometrically isomorphic designs.
@@
An examination of the results from the exchange algorithm indicated that designs with one centre point and any one of the nine points in the basis set occurring once and not twice are (DP)S-optimal. In terms of geometric isomorphism, consider now a design constructed directly from that of Gilmour and Trinca (2012) but with the the point now replicated twice and the point occurring once. For clarity, a stereogram of the design is shown in Figure 6(d), with the point occurring once indicated in red. It is clear that the design so constructed is not geometrically isomorphic to the design of Gilmour and Trinca (2012). However, by using arguments based on rotations and reflections, it is straightforward to show that the design represents a class of 24 isomorphic (DP)S-optimal designs. Furthermore, the same result holds for all designs constructed with one of the points in the set occurring once. There are therefore nine classes of designs, each comprising 24 geometrically isomorphic designs, and thus a total of designs.
@@
Consider now the (DP)S-optimal designs with two centre points generated from multiple runs of the exchange algorithm. A cursory examination of the results indicated that the points on the sphere of the design comprise two distinct points from the basis set chosen to occur once and the seven remaining points in replicated twice. The designs so defined were then constructed independently and shown to be (DP)S-optimal. It now follows from arguments similar to those presented for the (DP)S-optimal designs with one centre point that each of the designs so constructed represents a class of geometrically isomorphic designs, thereby yielding designs which are (DP)S-optimal and accounting for the designs generated by the exchange algorithm.
This example serves to illustrate the the importance of examining all optimal designs generated routinely by an exchange algorithm.
5. Discussion